# Private Online Community Detection for Censored Block Models

### v1

@@ D.1 The Synergy of Combining Multiple Adjacency Matrices
where $\ell(x)=\sup_{t\geq 0}-tx-\log\mathds{E}\left[e^{-tX_{1}}\right]$.
The term $\mathds{E}\left[e^{-tX_{1}}\right]$ is obtained as
$\mathds{E}\left[e^{-tX_{1}}\right]=1+p\left[e^{-t}(1-{\zeta})+e^{t}{\zeta}-1\right].$
$-x+\frac{{p}\left[e^{-t^{*}}(1-{\zeta})-e^{t^{*}}{\zeta}\right]}{1+{p}\left[e^{-t^{*}}(1-{\zeta})+e^{t^{*}}{\zeta}-1\right]}=0.$ (47)
For $x=o(1)$, the optimal parameter $t^{*}$ for the Chernoff bound is obtained as
$t^{*}$ $=\frac{1}{2}\log\frac{1-{\zeta}}{{\zeta}}+o(1).$ (48)
Plugging the value of $t^{*}$, it gives us the following:
$=-\frac{k_{n}}{2m}\log\frac{1-{\zeta}}{{\zeta}}+a\bigg{(}\sqrt{1-{\zeta}}-\sqrt{{\zeta}}\bigg{)}^{2}\frac{\log n}{n}+o(k_{n}/m)$
where the last equality follows due to the Taylor expansion of $\log(1-x)$ around $x=0$, and $p=a\log(n)/n$.
Thereby
$\exp(-m\ell(k_{n}/m))=\exp(-wa(\sqrt{1-{\zeta}}-\sqrt{{\zeta}})^{2}\log n+\frac{k_{n}}{2}\log\frac{1-{\zeta}}{{\zeta}}+o(k_{n}))$
choose $k_{n}=(1+o(1))\frac{\log n}{\log\log n}$ yields $\exp(-m\ell(k_{n}/m))=\exp(-wa(\sqrt{1-{\zeta}}-\sqrt{{\zeta}})^{2}\log n+\frac{k_{n}}{2}\log\frac{1-{\zeta}}{{\zeta}}+o(k_{n}))=n^{-wa(\sqrt{1-{\zeta}}-\sqrt{{\zeta}})^{2}+o(1)}$. Applying the union bound on $i$, we have with probability $1-n^{1-wa(\sqrt{1-{\zeta}}-\sqrt{{\zeta}})^{2}+o(1)}$,
$\min_{i}d_{i}^{\ast}\geq\frac{\log n}{\log\log n}.$
Second, we have
$\|\sum_{t=1}^{w}{\mathbf{A}}_{t}-\mathds{E}[\sum_{t=1}^{w}{\mathbf{A}}_{t}]\|\leq\sum_{t=1}^{w}\|{\mathbf{A}}_{t}-\mathds{E}[{\mathbf{A}}_{t}]\|$
by Theorem 9 in [17] we have $\operatorname{Pr}\{\|{\mathbf{A}}_{t}-\mathds{E}[{\mathbf{A}}_{t}]\|\leq c^{\prime}\sqrt{a\log n}\}\geq 1-n^{c}$, we have with probability at least $1-wn^{-c}$,
$\|\sum_{t=1}^{w}{\mathbf{A}}_{t}-\mathds{E}[\sum_{t=1}^{w}{\mathbf{A}}_{t}]\|\leq wc^{\prime}\sqrt{a\log n},$
for $p=a\frac{\log n}{n}$.
Combing these together we have with probability at least $1-n^{1-wa(\sqrt{1-{\zeta}}-\sqrt{{\zeta}})^{2}+o(1)}-wn^{-c}$, we have (46) $\geq\frac{\log n}{\log\log n}+w(1-2\zeta)p-wc^{\prime}\sqrt{a\log n}>0$ for $n$ sufficiently large. And therefore the model satisfies the exact recovery condition.
Appendix E Analysis of the Graph Perturbation Mechanism
In this section, we provide analysis for the Ternary Randomized Resonse Mechanism used in Section 4.1. In this mechanism, when $A_{i,j}=x$ where $x\in\{0,+1,-1\}$, the mechanism outputs $\tilde{A}_{i,j}=y$, where $y\in\{0,+1,-1\}$ with probability $q_{x}(y)\triangleq\operatorname{Pr}(\tilde{{\mathbf{A}}}_{i,j}=y|{\mathbf{A}}_{i,j}=x)$. We set $q_{x}(y)=c_{2}$, $\forall x\neq y$, and $q_{x}(y)=c_{1}=1-2c_{2}$, $\forall x=y$. We have the following:
$\tilde{p}_{+1}$ $\triangleq\operatorname{Pr}(\tilde{{\mathbf{A}}}_{i,j}=+1|\bm{\sigma}_{i}=\bm{\sigma}_{j})=c_{1}p(1-\zeta)+c_{2}p\zeta+c_{2}(1-p),$
$\tilde{p}_{-1}$ $\triangleq\operatorname{Pr}(\tilde{{\mathbf{A}}}_{i,j}=-1|\bm{\sigma}_{i}=\bm{\sigma}_{j})=c_{2}p(1-\zeta)+c_{1}p\zeta+c_{2}(1-p),$
$\tilde{p}_{0}$ $\triangleq\operatorname{Pr}(\tilde{{\mathbf{A}}}_{i,j}=0|\bm{\sigma}_{i}=\bm{\sigma}_{j})=c_{2}p(1-\zeta)+c_{2}p\zeta+c_{1}(1-p),$
$\tilde{q}_{+1}$ $\triangleq\operatorname{Pr}(\tilde{{\mathbf{A}}}_{i,j}=+1|\bm{\sigma}_{i}\neq\bm{\sigma}_{j})=c_{1}p\zeta+c_{2}p(1-\zeta)+c_{2}(1-p),$
$\tilde{q}_{-1}$ $\triangleq\operatorname{Pr}(\tilde{{\mathbf{A}}}_{i,j}=-1|\bm{\sigma}_{i}\neq\bm{\sigma}_{j})=c_{2}p\zeta+c_{1}p(1-\zeta)+c_{2}(1-p),$
$\tilde{q}_{0}$ $\triangleq\operatorname{Pr}(\tilde{{\mathbf{A}}}_{i,j}=0|\bm{\sigma}_{i}\neq\bm{\sigma}_{j})=c_{2}p(1-\zeta)+c_{2}p\zeta+c_{1}(1-p).$
Remark E.1 (CBM model after ternary-randomized perturbation).
We note that after the above ternary-randomized mechanism, the distribution of $\tilde{{\mathbf{A}}}$ is still a CBM model. More specifically, for the original CBM model ${\mathsf{CBM}}(\bm{\sigma},p,\zeta)$, the perturbed $\tilde{\mathbf{A}}$ follows the CBM model ${\mathsf{CBM}}(\bm{\sigma},\tilde{p},\tilde{\zeta})$ with the same community label vector $\bm{\sigma}$ and new parameters
$\tilde{p}=1-c_{1}(1-p)-c_{2}p=2c_{2}+p(1-3c_{2}),\,\tilde{\zeta}=\frac{c_{2}(1-p\zeta)+c_{1}p\zeta}{1-c_{1}(1-p)-c_{2}p}=\frac{c_{2}+p\zeta(1-3c_{2})}{2c_{2}+p(1-3c_{2})}.$ (49)
Remark E.2 (KL Divergence after Perturbation).
We first analyze the change in the KL divergence between the pre- and post-change distributions after the perturbation. The KL divergence for the raw data distribution (before perturbation) can be derived as (see Remark 5.2 for details),
$I(\bm{\sigma}^{\text{pre}};\bm{\sigma}^{\text{post}})=\left(\frac{1}{2}\log\frac{1-\zeta}{\zeta}\right)p(1-2\zeta)\left[{n\choose 2}-\sum_{i<j}\bm{\sigma}_{i}^{\text{pre}}\bm{\sigma}_{j}^{\text{pre}}\bm{\sigma}_{i}^{\text{post}}\bm{\sigma}_{j}^{\text{post}}\right].$
After the perturbation, we have similarly
$\tilde{I}(\bm{\sigma}^{\text{pre}};\bm{\sigma}^{\text{post}})=\left(\frac{1}{2}\log\frac{1-\tilde{\zeta}}{\tilde{\zeta}}\right)\tilde{p}(1-2\tilde{\zeta})\left[{n\choose 2}-\sum_{i<j}\bm{\sigma}_{i}^{\text{pre}}\bm{\sigma}_{j}^{\text{pre}}\bm{\sigma}_{i}^{\text{post}}\bm{\sigma}_{j}^{\text{post}}\right].$
Substitute the parameter $\tilde{p}$ and $\tilde{\zeta}$ in (49), and using the setting that $c_{1}=1-2c_{2}$, we have
$\left(\log\frac{1-\tilde{\zeta}}{\tilde{\zeta}}\right)\tilde{p}(1-2\tilde{\zeta})=\left(\log\frac{c_{2}+p(1-\zeta)(1-3c_{2})}{c_{2}+p\zeta(1-3c_{2})}\right)\cdot p(1-2\zeta)(1-3c_{2}).$ (50)
We note that $\log\frac{c_{2}+p(1-\zeta)(1-3c_{2})}{c_{2}+p\zeta(1-3c_{2})}$ is a positive and decreasing function of $c_{2}$ when $\zeta<\frac{1}{2}$ and a negative and increasing function when $\zeta>\frac{1}{2}$, the same is the function $p(1-2\zeta)(1-3c_{2})$. Moreover, when $c_{2}=0$ we have the KL divergence remaining unchanged. Combining these together, we have (50) is a nonnegative function and it is a decreasing function of $c_{2}$, with the largest value obtained when $c_{2}=0$ and the largest value equal to $I(\bm{\sigma}^{\text{pre}};\bm{\sigma}^{\text{post}})$. This demonstrates the tradeoff between detection and privacy. This being said, for a given privacy constraint, we may select the smallest value of $c_{2}$ value that satisfies this constraint; such choice ensures the best detection performance after the perturbation.
Note that, under $\epsilon$-edge LDP, we require $c_{2}/c_{1}\in[e^{-\epsilon},e^{\epsilon}]$, which yields the smallest possible value for $c_{2}$ being $c_{2}=\frac{1}{e^{\epsilon}+2}$, which is the value we choose in Section 4.1.
Next we derive the condition for exact recovery under the privacy mechanism.
Lemma E.1.
$\operatorname{Pr}\bigg{(}\sum_{i=1}^{m}X_{i}\leq k_{n}\bigg{)}\leq\exp(-m\ell(k_{n}/m)),$ (51)
@@ Appendix E Analysis of the Graph Perturbation Mechanism
$=-\frac{k_{n}}{2m}\log\frac{1-\tilde{\zeta}}{\tilde{\zeta}}-\log\left[1-\tilde{p}\bigg{(}\sqrt{1-\tilde{\zeta}}-\sqrt{\tilde{\zeta}}\bigg{)}^{2}\right]+o(k_{n}/m)$
=$\overset{(a)}{=}\tilde{p}\bigg{(}\sqrt{1-\tilde{\zeta}}-\sqrt{\tilde{\zeta}}\bigg{)}^{2}+o(\log(n)/n)$
$=\left[\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\times a+\frac{2}{e^{\epsilon}+2}\times\frac{n}{\log(n)}\right]$
×$\hskip 43.36243pt\times\bigg{(}\sqrt{\frac{1+(1-\zeta)p(e^{\epsilon}-1)}{2+p(e^{\epsilon}-1)}}-\sqrt{\frac{1+\zeta p(e^{\epsilon}-1)}{2+p(e^{\epsilon}-1)}}\bigg{)}^{2}\log(n)/n+o(\log(n)/n)$
$\geq\left[\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\times a+\frac{2}{e^{\epsilon}+2}\times\frac{n}{\log(n)}\right]$
×$\hskip 28.90755pt\times\bigg{(}\sqrt{\frac{(1-\zeta)p(e^{\epsilon}-1)}{2+p(e^{\epsilon}-1)}}-\sqrt{\frac{\zeta p(e^{\epsilon}-1)}{2+p(e^{\epsilon}-1)}}-\frac{1}{\sqrt{2+p(e^{\epsilon}-1)}}\bigg{)}^{2}\log(n)/n+o(\log(n)/n)$
$\geq\left[\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\times a+\frac{2}{e^{\epsilon}+2}\times\frac{n}{\log(n)}\right]$
×$\hskip 28.90755pt\times\frac{p(e^{\epsilon}-1)}{2+p(e^{\epsilon}-1)}\times\bigg{(}\sqrt{1-\zeta}-\sqrt{\zeta}-\frac{1}{\sqrt{p(e^{\epsilon}-1)}}\bigg{)}^{2}\log(n)/n+o(\log(n)/n)$
$\geq\left[\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\times a+\frac{2}{e^{\epsilon}+2}\times\frac{n}{\log(n)}\right]$
×$\hskip 28.90755pt\times\frac{p(e^{\epsilon}-1)}{2+p(e^{\epsilon}-1)}\times\bigg{(}\sqrt{1-\zeta}-\sqrt{\zeta}\bigg{)}^{2}\log(n)/n+o(\log(n)/n)$
@@
×$\hskip 28.90755pt\times\alpha\times\bigg{(}\sqrt{1-\zeta}-\sqrt{\zeta}\bigg{)}^{2}\log(n)/n+o(\log(n)/n),$ (53)
where step $(a)$ follows due to the Taylor expansion of $\log(1-x)$ around $x=0$, and $p=a\log(n)/n$ where the first term in step (a) behaves as $o(\log(n)/n)$. Step (b) holds true whenever $a>\frac{2(n^{3/2}-n)}{(n^{c}-1)\log(n)}$, where $\alpha=1-\frac{1}{\sqrt{n}}$, $\epsilon=c\log(n)$ and $c>0$. After applying the union bound, we get the following sufficient conditions for exact recovery:
$a\big{(}\sqrt{1-\zeta}-\sqrt{\zeta}\big{)}^{2}>\frac{e^{\epsilon}+1}{e^{\epsilon}-1}\times\left[1/\alpha-\frac{2}{e^{\epsilon}+2}\times\frac{n}{\log(n)}\times\big{(}\sqrt{1-\zeta}-\sqrt{\zeta}\big{)}^{2}\right].$
@@
$a\big{(}\sqrt{1-\zeta}-\sqrt{\zeta}\big{)}^{2}>\frac{1}{\alpha}\times\left(\frac{e^{\epsilon}+1}{e^{\epsilon}-1}\right).$
Remark E.3.
In conclusion, the analysis of the lower bounds on the parameter $a$ for different formulations of $\alpha$, specifically $\alpha=1-\frac{1}{\sqrt{n}}$ and $\alpha=1-\frac{1}{n}$, reveals significant variations in the required values of $a$ for a given $c$. Notably, for a fixed $c$, the choice of $\alpha=1-\frac{1}{\sqrt{n}}$ generally results in a tighter (or lower) lower bound for $a$ across a range of $n$ values. This tighter bound is preferable as it allows for more flexibility and potentially less stringent requirements on the parameter $a$ while maintaining the desired privacy level defined by $\epsilon=c\log(n)$. Thus, in scenarios where achieving an optimal balance between privacy and parameter constraints is crucial, opting for $\alpha=1-\frac{1}{\sqrt{n}}$ would be a more strategic choice.
Appendix F Fundamental Limits of Private Community Recovery
@@
Lemma F.1.
(Packing lower bound for graph datasets) Let $\mathbf{A}_{1},\mathbf{A}_{2},\cdots\mathbf{A}_{m}\in\{0,-1,1\}^{n\times n}$ be a set of $m$ graph datasets . For any pair of datasets $\mathbf{A}_{i}$ and $\mathbf{A}_{j}$, we have $\operatorname{Ham}(\mathbf{A}_{i},\mathbf{A}_{j})\leq d$. Let $\mathcal{S}_{1},\mathcal{S}_{2},\cdots,\mathcal{S}_{m}$ be a set of $m$ disjoint subsets. If there exists an $\epsilon$-DP mechanism $\mathcal{M}:\mathcal{G}\rightarrow\mathcal{S}$ such that $\operatorname{Pr}(\mathcal{M}(G_{i})\in\mathcal{S}_{i})\geq 9/10$ for every $i\in[m]$, then $\epsilon=\Omega(\log(m)/d)$.
Remark F.1.
Note that $d={n\choose 2}$ is the worst case upper bound on the Hamming distance between any two graphs (i.e., adjacency matrices). We next derive similar type of lower bounds that are stronger by replacing the worst-case distance with the expected hamming distance between any two random graphs generated from a joint distribution.
F.1 Converse for $\epsilon$-edge DP Community Recovery Algorithms
We establish a rigorous lower bound for all differentially private community recovery algorithms operating on graphs generated from the Constrained Block Models (CBMs). Our methodology closely follows the frameworks outlined in [33, 6], focusing on the notion of edge DP. Furthermore, we precisely define the classification error rate as
${\operatorname{err}}\operatorname{rate}(\hat{\bm{\sigma}}(\mathbf{A}),{\bm{\sigma}^{*}})$ $=\frac{1}{n}\times\min\{\operatorname{Ham}(\hat{\bm{\sigma}}(\mathbf{A}),{\bm{\sigma}^{*}}),\operatorname{Ham}(-\hat{\bm{\sigma}}(\mathbf{A}),{\bm{\sigma}^{*}})\}.$ (54)
Let us consider a series of pairwise disjoint sets $\mathcal{S}_{i}$ for $i=1,2,\ldots,m$. Each set $\mathcal{S}_{i}$ contains vectors $\bm{u}\in\{\pm 1\}^{n}$, where $n$ is the vector dimension. A vector $\bm{u}$ is included in $\mathcal{S}_{i}$ if the error rate $\operatorname{err}\operatorname{rate}(\bm{u},\bm{\sigma})$ with a fixed vector $\bm{\sigma}$ does not exceed the threshold $\beta$. This is formally expressed as:
$\mathcal{S}_{i}=\{\bm{u}\in\{\pm 1\}^{n}:\operatorname{err}\operatorname{rate}(\bm{u},\bm{\sigma})\leq\beta\},\quad i=1,2,\ldots,m,$ (55)
@@ F.1 Converse for $\epsilon$-edge DP Community Recovery Algorithms
Further, for each index $i$, let $P_{i}$ be the distribution over $n$-vertex graphs generated by an CBM with parameters $\zeta$, $p$ and $\bm{\sigma}^{i}$, in short $\operatorname{CBM}(\zeta,p,\sigma^{i})$. We next derive the necessary conditions for
@@
as a function of the CBM parameters and the privacy budget. Note that the randomness here is taken over the randomness of graph $G$ that is generated from $\operatorname{CBM}(\zeta,p,\sigma^{i})$. Without loss of generality, let us consider a graph $\mathbf{A}\sim\operatorname{CBM}(\zeta,p,\sigma^{1})$ that is generated from the ground truth labeling vector $\bm{\sigma}^{*}=\bm{\sigma}^{1}$. Further, for this case, we want to derive the necessary conditions for any $\epsilon$-edge DP recovery algorithms that
@@
where $\mathbf{A}\sim\operatorname{CBM}(\zeta,p,\sigma^{1})$.
@@
$\mathds{E}_{\mathbf{A},\mathbf{A}^{\prime}\sim\Pi(\mathbf{A},\mathbf{A}^{\prime})}\left[\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A})\in\mathcal{S}_{i})\right]$ $\leq\mathds{E}_{\mathbf{A},\mathbf{A}^{\prime}\sim\Pi(\mathbf{A},\mathbf{A}^{\prime})}\left[e^{\epsilon\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime})}\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A^{\prime}})\in\mathcal{S}_{i})\right]$
$\Rightarrow\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A})\in\mathcal{S}_{i})$ $\leq\mathds{E}_{\mathbf{A},\mathbf{A}^{\prime}\sim\Pi(\mathbf{A},\mathbf{A}^{\prime})}\left[e^{\epsilon\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime})}\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A^{\prime}})\in\mathcal{S}_{i})\right]$
@@
×$\hskip 93.95122pt\times\left(\mathds{E}_{\mathbf{A},\mathbf{A}^{\prime}\sim\Pi(\mathbf{A},\mathbf{A}^{\prime})}\left[\operatorname{Pr}^{2}(\hat{\bm{\sigma}}(\mathbf{A^{\prime}})\in\mathcal{S}_{i})\right]\right)^{1/2}$
$\leq\left(\mathds{E}_{\mathbf{A},\mathbf{A}^{\prime}\sim\Pi(\mathbf{A},\mathbf{A}^{\prime})}\left[e^{2\epsilon\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime})}\right]\right)^{1/2}\times\left(\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A^{\prime}})\in\mathcal{S}_{i})\right)^{1/2},$
$\Rightarrow 1-\eta$ $\overset{(b)}{\leq}\left(\mathds{E}_{\mathbf{A},\mathbf{A}^{\prime}\sim\Pi(\mathbf{A},\mathbf{A}^{\prime})}\left[e^{2\epsilon\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime})}\right]\right)^{1/2}\times\left(\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A^{\prime}})\in\mathcal{S}_{i})\right)^{1/2},$
$\Rightarrow(1-\eta)^{2}$ $\leq\left(\mathds{E}_{\mathbf{A},\mathbf{A}^{\prime}\sim\Pi(\mathbf{A},\mathbf{A}^{\prime})}\left[e^{2\epsilon\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime})}\right]\right)\times\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A^{\prime}})\in\mathcal{S}_{i}),$ (59)
@@
F.2 Characterizing the randomness of the Hamming distance between two random adjacency matrices
@@
$M_{\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime})}(2\epsilon)$ $=\left(M_{\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime}):\text{same dist.}}(2\epsilon)\right)^{(n-\operatorname{Ham}(\bm{\sigma},\bm{\sigma}^{\prime}))}\times\left(M_{\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime}):\text{different dist.}}(2\epsilon)\right)^{\operatorname{Ham}(\bm{\sigma},\bm{\sigma}^{\prime})},$ (60)
@@
$M_{\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime})}(2\epsilon)$ $\leq\left(M_{\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime}):\text{same dist.}}(2\epsilon)\right)^{(n-\operatorname{Ham}(\bm{\sigma},\bm{\sigma}^{\prime}))\cdot\operatorname{Ham}(\bm{\sigma},\bm{\sigma}^{\prime})}$
$=\left(-2p^{2}\zeta(\zeta-1)+(p-1)^{2}+\left(2p^{2}\zeta(\zeta-1)-(p-1)^{2}+1\right)e^{2\epsilon}\right)^{(n-\operatorname{Ham}(\bm{\sigma},\bm{\sigma}^{\prime}))\cdot\operatorname{Ham}(\bm{\sigma},\bm{\sigma}^{\prime})}$
@@
F.3 Lower bounding the packing number $m$ with respect to $\operatorname{err}\operatorname{rate}$:
Building upon the framework established in [6], we can readily demonstrate that

### v2

@@ Appendix C Proofs for Section IV-A
$=-\frac{k_{n}}{2m}\log\frac{1-\tilde{\zeta}}{\tilde{\zeta}}-\log\left[1-\tilde{p}\bigg(\sqrt{1-\tilde{\zeta}}-\sqrt{\tilde{\zeta}}\bigg)^{2}\right]$
$\qquad+o(k_{n}/m)$
$\overset{(a)}{=}\tilde{p}\bigg(\sqrt{1-\tilde{\zeta}}-\sqrt{\tilde{\zeta}}\bigg)^{2}+o(\log(n)/n),$
where step $(a)$ follows due to the Taylor expansion of $\log(1-x)$ around $x=0$, and $p=a\log(n)/n$ where the first term in step (a) behaves as $o(\log(n)/n)$. Substitute $\tilde{p}=\frac{2+p(e^{\epsilon}-1)}{e^{\epsilon}+2}$ and $\tilde{\zeta}=\frac{1+p\zeta(e^{\epsilon}-1)}{2+p(e^{\epsilon}-1)}$, we obtain
$\ell(k_{n}/m)\geq\left[\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\times a+\frac{2}{e^{\epsilon}+2}\times\frac{n}{\log(n)}\right]$
$\hskip 36.135pt\times\frac{p(e^{\epsilon}-1)}{2+p(e^{\epsilon}-1)}\times\bigg(\sqrt{1-\zeta}-\sqrt{\zeta}\bigg)^{2}\log(n)/n$
$\hskip 36.135pt+o(\log(n)/n)$
@@
$\hskip 14.45377pt\times\alpha\times\bigg(\sqrt{1-\zeta}-\sqrt{\zeta}\bigg)^{2}\log(n)/n+o(\log(n)/n),$
where step (b) holds true whenever $a>\frac{2(n^{3/2}-n)}{(n^{c}-1)\log(n)}$, where $\alpha=1-\frac{1}{\sqrt{n}}$, $\epsilon=c\log(n)$ and $c>0$. After applying the union bound, we get the following sufficient conditions for exact recovery is:
$a\big(\sqrt{1-\zeta}-\sqrt{\zeta}\big)^{2}>\frac{e^{\epsilon}+1}{e^{\epsilon}-1}$
$\hskip 21.68121pt\times\bigg[1/\alpha-\frac{2}{e^{\epsilon}+2}\times\frac{n}{\log(n)}\times\big(\sqrt{1-\zeta}-\sqrt{\zeta}\big)^{2}\bigg].$ (32)
@@
$a\big(\sqrt{1-\zeta}-\sqrt{\zeta}\big)^{2}>\frac{1}{\alpha}\times\left(\frac{e^{\epsilon}+1}{e^{\epsilon}-1}\right),$
thereby we prove (12) in Theorem IV.1.
C-B Proof of Theorem IV.4
@@
We establish a rigorous lower bound for all differentially private community recovery algorithms operating on graphs generated from the Censored Block Models (CBMs). Our methodology closely follows the frameworks outlined in [37, 7], focusing on the notion of edge DP. Furthermore, we precisely define the classification error rate as
${\operatorname{err}}\operatorname{rate}(\hat{\bm{\sigma}}(\mathbf{A}),{\bm{\sigma}^{*}})$
$\hskip 7.22743pt=\frac{1}{n}\times\min\{\operatorname{Ham}(\hat{\bm{\sigma}}(\mathbf{A}),{\bm{\sigma}^{*}}),\operatorname{Ham}(-\hat{\bm{\sigma}}(\mathbf{A}),{\bm{\sigma}^{*}})\}.$ (33)
Let us consider a series of pairwise disjoint sets $\mathcal{S}_{i},i\in[m]$. Each set $\mathcal{S}_{i}$ contains vectors $\bm{u}\in\{\pm 1\}^{n}$, where $n$ is the vector dimension. A vector $\bm{u}$ is included in $\mathcal{S}_{i}$ if the error rate $\operatorname{err}\operatorname{rate}(\bm{u},\bm{\sigma}^{i})$ with a fixed vector $\bm{\sigma}^{i}$ does not exceed the threshold $\beta$. This is formally expressed as:
$\mathcal{S}_{i}=\{\bm{u}\in\{\pm 1\}^{n}:\operatorname{err}\operatorname{rate}(\bm{u},\bm{\sigma}^{i})\leq\beta\},\qquad i=1,2,\ldots,m,$ (34)
@@
We next derive the necessary conditions for
@@
as a function of the CBM parameters and the privacy budget. Note that the randomness here is taken over the randomness of graph $G$ that is generated from $\operatorname{CBM}(\bm{\sigma}^{i},p,\zeta)$. Without loss of generality, let us consider a graph $\mathbf{A}\sim\operatorname{CBM}(\bm{\sigma}^{1},p,\zeta)$ that is generated from the ground truth labeling vector $\bm{\sigma}^{*}=\bm{\sigma}^{1}$. Further, for this case, we want to derive the necessary conditions for any $\epsilon$-edge DP recovery algorithms that
@@
where $\mathbf{A}\sim\operatorname{CBM}(\bm{\sigma}^{1},p,\zeta)$.
@@
$\mathds{E}_{\mathbf{A},\mathbf{A}^{\prime}\sim\Pi(\mathbf{A},\mathbf{A}^{\prime})}\left[\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A})\in\mathcal{S}_{i})\right]$
$\hskip 14.45377pt\leq\mathds{E}_{\mathbf{A},\mathbf{A}^{\prime}\sim\Pi(\mathbf{A},\mathbf{A}^{\prime})}\left[e^{\epsilon\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime})}\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A^{\prime}})\in\mathcal{S}_{i})\right]$
$\Rightarrow\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A})\in\mathcal{S}_{i})$
$\hskip 14.45377pt\leq\mathds{E}_{\mathbf{A},\mathbf{A}^{\prime}\sim\Pi(\mathbf{A},\mathbf{A}^{\prime})}\left[e^{\epsilon\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime})}\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A^{\prime}})\in\mathcal{S}_{i})\right]$
@@
$\hskip 72.26999pt\times\left(\mathds{E}_{\mathbf{A},\mathbf{A}^{\prime}\sim\Pi(\mathbf{A},\mathbf{A}^{\prime})}\left[\operatorname{Pr}^{2}(\hat{\bm{\sigma}}(\mathbf{A^{\prime}})\in\mathcal{S}_{i})\right]\right)^{1/2}$
$\leq\left(\mathds{E}_{\mathbf{A},\mathbf{A}^{\prime}\sim\Pi(\mathbf{A},\mathbf{A}^{\prime})}\left[e^{2\epsilon\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime})}\right]\right)^{1/2}$
$\hskip 72.26999pt\times\left(\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A^{\prime}})\in\mathcal{S}_{i})\right)^{1/2},$
$\Rightarrow 1-\eta\overset{(b)}{\leq}\left(\mathds{E}_{\mathbf{A},\mathbf{A}^{\prime}\sim\Pi(\mathbf{A},\mathbf{A}^{\prime})}\left[e^{2\epsilon\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime})}\right]\right)^{1/2}$
$\hskip 72.26999pt\times\left(\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A^{\prime}})\in\mathcal{S}_{i})\right)^{1/2},$
$\Rightarrow(1-\eta)^{2}\leq\left(\mathds{E}_{\mathbf{A},\mathbf{A}^{\prime}\sim\Pi(\mathbf{A},\mathbf{A}^{\prime})}\left[e^{2\epsilon\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime})}\right]\right)$
$\hskip 72.26999pt\times\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A^{\prime}})\in\mathcal{S}_{i}),$ (38)
@@
$M_{\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime})}(2\epsilon)$ $=\left(M_{\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime}):\text{same dist.}}(2\epsilon)\right)^{(n-\operatorname{Ham}(\bm{\sigma},\bm{\sigma}^{\prime}))}$
$\hskip 7.22743pt\times\left(M_{\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime}):\text{different dist.}}(2\epsilon)\right)^{\operatorname{Ham}(\bm{\sigma},\bm{\sigma}^{\prime})},$ (39)
@@
$M_{\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime})}(2\epsilon)$
$\leq\left(M_{\operatorname{Ham}(\mathbf{A},\mathbf{A}^{\prime}):\text{same dist.}}(2\epsilon)\right)^{(n-\operatorname{Ham}(\bm{\sigma},\bm{\sigma}^{\prime}))\cdot\operatorname{Ham}(\bm{\sigma},\bm{\sigma}^{\prime})}$
@@
Finally, we lower bound the packing number $m$ with respect to $\operatorname{err}\operatorname{rate}$. Building upon the framework established in [7], we can readily demonstrate that